OpenStudy (anonymous):

Please check my answer? 5^4√7 - 8^3√7 - 2^4√7 - ^3√7 3^4√7-9^3√7 -6√14 -6√7 3^4√14 - 9^3√14 My answer is C

OpenStudy (anonymous):

Why you have used power sign there ^ ??

OpenStudy (anonymous):

That's how it's written :)

OpenStudy (anonymous):

Wait my internet is very slow now..

OpenStudy (anonymous):

Is the file attached by you is your final question??

OpenStudy (anonymous):

It's okay :) My internet gets slow too, plus I have a Dell from hell

OpenStudy (anonymous):

Is this your question?? \[\large 5 \sqrt[4]{7} - 8 \sqrt[3]{7} - 2 \sqrt[4]{7} - \sqrt[3]{7}\]

OpenStudy (anonymous):

Hey Dell is a good company no??

OpenStudy (anonymous):

It's a terrible company for me :( and yes!

OpenStudy (anonymous):

Then C looks not nice to me..

OpenStudy (anonymous):

What looks nice to you then/ Can you explain why? :)

OpenStudy (anonymous):

See you can add or subtract the like terms only in Mathematics.. \(\sqrt[3]{7}\) and \(\sqrt[4]{7}\) are not like terms you cannot add or subtract them.. Okay??

OpenStudy (anonymous):

You can only add or subtract : \(4 \sqrt[3]{7}\) and \(3 \sqrt[3]{7}\)

OpenStudy (anonymous):

The nth power on root must be same. If it is 3 then in both the terms it should be 3..

OpenStudy (anonymous):

Now can you tell what things in your question can you add or subtract??

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The exponents?

OpenStudy (anonymous):

3x3?

OpenStudy (anonymous):

See don't get confuse : Say you can solve this by substitute method it will become easy for you then: Put : \(\sqrt[3]{7}\) = x and \(\sqrt[4]{7}\) = y : Now your question has become: \[5y - 8x - 2y- x\] Can you solve this or not ??

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Tell me : \(5y - 2y = ??\)

OpenStudy (anonymous):

3y

OpenStudy (anonymous):

Yes, similar way : \(-8x - x = ??\)

OpenStudy (anonymous):

I don't know about that :/

OpenStudy (anonymous):

- and - in case of addition is : (-) (-) becomes +.. but sign does not change: So: -8x - x = -9x If both are negative, then value gets added but sign remains negative..

OpenStudy (anonymous):

Okay :)